use elimination to slove each system of equations
3x+1=-7y
6x+7y=0

Answer by Alan3354(69443)

use elimination to slove each system of equations
3x+1=-7y
6x+7y=0
=====================
3x+7y=-1
6x+7y=0
--------- Subtract